Solve the system by graphing.
x – 2y = 8
x + y = -1



Answer by daniellagirl89(3)

x – 2y = 8
x + y = -1
Multiply the second equation by 2 to get both y's opposite of eachother
2(x + y) = -1(2)
2x + 2y = -2
Now plug in your first equation
x - 2y = 8
2x + 2y = -2
3x = 6
x = 2